Each of two small spheres is charged positively, the total charge being 52.6 wC. Each sphere is repelled from the other with a force of 1.19 N when the spheres are 1.94 m apart. Calculate the charge on each sphere. 


1
Expert's answer
2022-02-02T17:34:49-0500

"F=\frac{kq^2}{r^2},\implies"

"q=r\sqrt{\frac Fk}=22.3~\mu C."